OpenStudy (anonymous):

3a(a^2+5)-11 = a(3a^2+4) Solve for a Please include steps, thanks.

OpenStudy (anonymous):

\[3a(a^2+5)-11=a(3a^2+4)\]

OpenStudy (anonymous):

\[3a^3+15a-11=3a^3+4a\]

OpenStudy (anonymous):

then get rid of the \[3a^3\] to get \[15a-11=4a\] \[11a=11\] \[a=1\]

OpenStudy (anonymous):

Thanks